iosangel:

solve for x 3^7x=81^x+9

iosangel:



iosangel:

Can i use log to solve this

Nnesha:

yes

surjithayer:

any idea?

surjithayer:

no,need to take log. \[3^{7x}=81^{x+9}\] \[3^{7x}=(3^4)^{(x+9)}\] \[3^{7x}=3^{4(x+9)}\] \[3^{7x}=3^{(4x+36)}\] 7x=4x+36 x=?

iosangel:

12